# Wasserstein Proximal Coordinate Gradient Algorithms

### v1

@@ Appendix A Proof of Main Results
where $C$ is a constant varying from line to line. Here, (i) is by the inequality (24) in both the parallel and the sequential update schemes; (ii) is by Lemma A.1 in the parallel update scheme and Lemma A.3 in the sequential update scheme.
@@
where the last inequality is due the non-increasing property of $\mathcal{F}$ derived by Lemma A.1 and Lemma A.3. This implies
@@
Similar to the proof of Theorem 4.8, we consider the decomposition
@@
By (25) and Lemma A.5, the first term is bounded by
@@
A.5 Inexact WPCG-P with $\lambda>0$: proof of Theorem 4.13
The proof is similar to the exact WPCG, but we need to take the cumulative error into consideration. Some analogous lemmas as in the proof of Theorem 4.2 are required, the proofs of which are postponed to Appendix B.
Lemma A.7 (sufficient decrease for inexact WPCG-P).
@@
Lemma A.8 (bound of functional value for inexact WPCG-P).
@@
Here (i) is by Lemma A.8, and (ii) is by Lemma A.7. Reorganizing the inequality yields
@@
Applying Lemma C.3 with
@@
A.6 Inexact WPCG-P with $\lambda>0$: proof of Theorem 4.15
@@
$\leq\sqrt{3}D_{\mathcal{X}}\bigg{[}\frac{\|\varepsilon^{k}\|}{\tau}+\sqrt{\Big{(}L^{2}(m-1)+\frac{1}{\tau^{2}}\Big{)}\Big{(}\frac{1}{2\tau}-L\sqrt{m-1}\Big{)}^{-1}\Big{(}\mathcal{F}(\widetilde{\rho}^{k-1})-\mathcal{F}(\widetilde{\rho}^{k})+\frac{\|\varepsilon^{k}\|^{2}}{2\tau}\Big{)}}\bigg{]}.$
Here, we apply Lemma A.7 and the fact that $\operatorname{W}_{2}(\widetilde{\rho}^{k},\rho^{\ast})\leq D_{\mathcal{X}}$. If $\mathcal{F}(\widetilde{\rho}^{k})-\mathcal{F}(\rho^{\ast})\geq\sqrt{3}D_{\mathcal{X}}\frac{\|\varepsilon^{k}\|}{\tau}$, we have
@@
Applying Lemma C.4 with
@@ Appendix B Proof of Lemmas in the Main Theorems
To prove the two steps mentioned in the sketch at the beginning of Appendix A, the most important result is the following first-order optimality condition (FOC). The proof, which will be postponed to Appendix C.2, is similar to Proposition 8.7 in [42] for $h(x)=x\log x$ and $W=0$.
Lemma B.1 (FOC).
@@
By Lemma B.1, we have $T_{\rho_{j}^{k+1}}^{\rho_{j}^{k}}-\textrm{Id}=\tau\nabla V_{j}^{k}+\tau\nabla h_{j}^{\prime}(\rho_{j}^{k+1})+\tau\nabla_{\operatorname{W}_{2}}\mathcal{W}_{j}(\rho_{j}^{k+1})$. This implies
@@
$=\sum_{j=1}^{m}\int_{\mathcal{X}_{j}}\big{\langle}\nabla V_{j}^{k+1},T_{\rho_{j}^{k+1}}^{\rho_{j}^{k}}-\textrm{Id}\big{\rangle}+\bigg{\langle}\frac{T_{\rho_{j}^{k+1}}^{\rho_{j}^{k}}-\textrm{Id}}{\tau}-\big{(}\nabla V_{j}^{k}+\nabla h_{j}^{\prime}(\rho_{j}^{k+1})+\nabla_{\operatorname{W}_{2}}\mathcal{W}_{j}(\rho_{j}^{k+1})\big{)},T_{\rho_{j}^{k+1}}^{\rho_{j}^{k}}-\textrm{Id}\bigg{\rangle}\,{\rm d}\rho_{j}^{k+1}$
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\int_{\mathcal{X}_{j}}\Big{\langle}\nabla_{\operatorname{W}_{2}}\mathcal{W}_{j}(\rho_{j}^{k+1}),T_{\rho_{j}^{k+1}}^{\rho_{j}^{k}}-\textrm{Id}\Big{\rangle}\,{\rm d}\rho_{j}^{k+1}+\frac{1}{\tau}\operatorname{W}_{2}^{2}(\rho_{j}^{k+1},\rho_{j}^{k})\bigg{]}$
$\stackrel{{\scriptstyle(\textrm{i})}}{{\geq}}\sum_{j=1}^{m}\bigg{[}\int_{\mathcal{X}_{j}}\Big{\langle}\nabla V_{j}^{k+1}-\nabla V_{j}^{k},T_{\rho_{j}^{k+1}}^{\rho_{j}^{k}}-\textrm{Id}\Big{\rangle}\,{\rm d}\rho_{j}^{k+1}-\Big{(}\mathcal{H}_{j}(\rho_{j}^{k})-\mathcal{H}_{j}(\rho_{j}^{k+1})\Big{)}-\Big{(}\mathcal{W}_{j}(\rho_{j}^{k})-\mathcal{W}_{j}(\rho_{j}^{k+1})\Big{)}\bigg{]}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\frac{1}{\tau}\operatorname{W}_{2}^{2}(\rho^{k+1},\rho^{k})$
$\stackrel{{\scriptstyle(\textrm{ii})}}{{\geq}}-\sum_{j=1}^{m}\operatorname{W}_{2}(q_{j}^{k+1},q_{j}^{k})\cdot\Big{\|}\nabla V_{j}^{k+1}-\nabla V_{j}^{k}\Big{\|}_{L^{2}(\rho_{j}^{k+1};\mathcal{X}_{j})}-\Big{(}\mathcal{H}(\rho^{k})-\mathcal{H}(\rho^{k+1})\Big{)}-\Big{(}\mathcal{W}(\rho^{k})-\mathcal{W}(\rho^{k+1})\Big{)}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\frac{1}{\tau}\operatorname{W}_{2}^{2}(\rho^{k+1},\rho^{k})$
$\stackrel{{\scriptstyle(\textrm{iii})}}{{\geq}}-\sum_{j=1}^{m}\operatorname{W}_{2}(\rho_{j}^{k+1},\rho_{j}^{k})\cdot L\operatorname{W}_{2}(\rho_{-j}^{k+1},\rho_{-j}^{k})-\Big{(}\mathcal{H}(\rho^{k})-\mathcal{H}(\rho^{k+1})\Big{)}-\Big{(}\mathcal{W}(\rho^{k})-\mathcal{W}(\rho^{k+1})\Big{)}+\frac{1}{\tau}\operatorname{W}_{2}^{2}(\rho^{k+1},\rho^{k})$
$\stackrel{{\scriptstyle\textrm{(iv)}}}{{\geq}}-L\sum_{j=1}^{m}\bigg{[}\frac{\sqrt{m-1}}{2}\operatorname{W}_{2}^{2}(\rho_{j}^{k+1},\rho_{j}^{k})+\frac{1}{2\sqrt{m-1}}\operatorname{W}_{2}^{2}(\rho_{-j}^{k+1},\rho_{-j}^{k})\bigg{]}-\Big{(}\mathcal{H}(\rho^{k})-\mathcal{H}(\rho^{k+1})\Big{)}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\Big{(}\mathcal{W}(\rho^{k})-\mathcal{W}(\rho^{k+1})\Big{)}+\frac{1}{\tau}\operatorname{W}_{2}^{2}(\rho^{k+1},\rho^{k})$
@@
Here, (i) is due to the fact that both $\mathcal{H}_{j}$ and $\mathcal{W}_{j}$ are convex along geodesics; (ii) is by Cauchy–Schwarz inequality; (iii) is by Lemma C.1; (iv) is by AM-GM inequality.
@@
$\geq m\mathcal{V}(\rho^{k+1})+(m-1)\Big{[}\big{(}\tau^{-1}-L\sqrt{m-1}\big{)}\operatorname{W}_{2}^{2}(\rho^{k+1},\rho^{k})-\Big{(}\mathcal{H}(\rho^{k})-\mathcal{H}(\rho^{k+1})\Big{)}-\Big{(}\mathcal{W}(\rho^{k})-\mathcal{W}(\rho^{k+1})\Big{)}\Big{]}.$
@@
Here, both (i) and (ii) are due to Cauchy–Schwarz inequality, and (iii) is by Lemma B.2. By Assumption D ($\lambda$-QG condition), we have
@@
B.3 Proof of Lemma A.3
Proof.
By the definition of $\rho_{j}^{k}$, we have
@@
B.4 Proof of Lemma A.4
Proof.
Let $Y_{j}=\sum_{l=0}^{M-1}I\{j_{l}=j\}$ and $A_{j}=\{Y_{j}=0\}\cup\{Y_{j}>T\}$ be a event. Then
@@
Applying Chernoff’s inequality to Bernoulli distribution (Theorem 2.3.1 in[47]) yields
@@
we have $2m^{1-R}=2\delta$. ∎
B.5 Proof of Lemma A.5
Proof.
By the definition of $\rho_{j_{l}}^{k,l+1}$, we have
@@
Here, both (i) and (ii) are due to $\rho_{-j_{l}}^{k,l+1}=\rho_{-j_{l}}^{k,l}$. ∎
B.6 Proof of Lemma A.6
Proof.
Similar to the argument in Lemma A.2, we have
@@
$\sum_{j=1}^{m}\mathcal{V}(\widetilde{\rho}_{j}^{k+1}\otimes\widetilde{\rho}_{-j}^{k})\leq m\mathcal{V}(\widetilde{\rho}^{k})+\mathcal{H}(\widetilde{\rho}^{k})+\mathcal{W}(\widetilde{\rho}^{k})-\mathcal{H}(\widetilde{\rho}^{k+1})-\mathcal{W}(\widetilde{\rho}^{k})-\frac{1}{2\tau}\operatorname{W}_{2}^{2}(\widetilde{\rho}^{k+1},\widetilde{\rho}^{k})+\frac{\|\varepsilon^{k+1}\|^{2}}{2\tau}.$
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\Big{\langle}\nabla V_{j}^{k+1},T_{\widetilde{\rho}_{j}^{k+1}}^{\widetilde{\rho}_{j}^{k}}-\textrm{Id}\Big{\rangle}\,{\rm d}\widetilde{\rho}_{j}^{k+1}$
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\int_{\mathcal{X}_{j}}\Big{\langle}\nabla_{\operatorname{W}_{2}}\mathcal{W}_{j}(\widetilde{\rho}_{j}^{k+1}),T_{\widetilde{\rho}_{j}^{k+1}}^{\widetilde{\rho}_{j}^{k}}-\textrm{Id}\Big{\rangle}\,{\rm d}\widetilde{\rho}_{j}^{k+1}+\frac{1}{\tau}\operatorname{W}_{2}^{2}(\widetilde{\rho}_{j}^{k+1},\widetilde{\rho}_{j}^{k})$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad-\frac{1}{\tau}\int_{\mathcal{X}_{j}}\Big{\langle}\eta_{j}^{k+1},T_{\widetilde{\rho}_{j}^{k+1}}^{\widetilde{\rho}_{j}^{k}}-\textrm{Id}\Big{\rangle}\,{\rm d}\widetilde{\rho}_{j}^{k+1}\bigg{]}$
@@
$\qquad\qquad\qquad\qquad\quad-\int_{\mathcal{X}_{j}}\Big{\langle}\nabla_{\operatorname{W}_{2}}\mathcal{W}_{j}(\widetilde{\rho}_{j}^{k+1}),T_{\widetilde{\rho}_{j}^{k+1}}^{\widetilde{\rho}_{j}^{k}}-\textrm{Id}\Big{\rangle}\,{\rm d}\widetilde{\rho}_{j}^{k+1}+\frac{1}{2\tau}\operatorname{W}_{2}^{2}(\widetilde{\rho}_{j}^{k+1},\widetilde{\rho}_{j}^{k})-\frac{(\varepsilon_{j}^{k+1})^{2}}{2\tau}\bigg{]}$
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\big{[}\mathcal{W}(\widetilde{\rho}^{k})-\mathcal{W}(\widetilde{\rho}^{k+1})\big{]}+\frac{1}{2\tau}\operatorname{W}_{2}^{2}(\widetilde{\rho}^{k+1},\widetilde{\rho}^{k})-\frac{\|\varepsilon^{k+1}\|^{2}}{2\tau}$
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\big{[}\mathcal{W}(\widetilde{\rho}^{k})-\mathcal{W}(\widetilde{\rho}^{k+1})\big{]}+\frac{1}{2\tau}\operatorname{W}_{2}^{2}(\widetilde{\rho}^{k+1},\widetilde{\rho}^{k})-\frac{\|\varepsilon^{k+1}\|^{2}}{2\tau}$
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\big{[}\mathcal{W}(\widetilde{\rho}^{k})-\mathcal{W}(\widetilde{\rho}^{k+1})\big{]}+\frac{1}{2\tau}\operatorname{W}_{2}^{2}(\widetilde{\rho}^{k+1},\widetilde{\rho}^{k})-\frac{\|\varepsilon^{k+1}\|^{2}}{2\tau}$

### v2

@@ A.4 Case $\lambda=0$: Proof of Theorem 18
where $C$ is a constant varying from line to line. Here, (i) is by the inequality (24) in both the parallel and the sequential update schemes; (ii) is by Lemma 1 in the parallel update scheme and Lemma 3 in the sequential update scheme.
@@
where the last inequality is due the non-increasing property of $\mathcal{F}$ derived by Lemma 1 and Lemma 3. This implies
@@
Similar to the proof of Theorem 15, we consider the decomposition
@@
By (25) and Lemma 5, the first term is bounded by
@@
A.5 Inexact WPCG-P with $\lambda>0$: Proof of Theorem 20
The proof is similar to the exact WPCG, but we need to take the cumulative error into consideration. Some analogous lemmas as in the proof of Theorem 9 are required, the proofs of which are postponed to Appendix B.
Lemma 7 (sufficient decrease for inexact WPCG-P).
@@ A.5 Inexact WPCG-P with $\lambda>0$: Proof of Theorem 20
Lemma 8 (bound of functional value for inexact WPCG-P).
@@
Here (i) is by Lemma 8, and (ii) is by Lemma 7. Reorganizing the inequality yields
@@
Applying Lemma 14 with
@@
A.6 Inexact WPCG-P with $\lambda>0$: Proof of Theorem 22
@@
$\leq\sqrt{3}D_{\mathcal{X}}\bigg{[}\frac{\|\varepsilon^{k}\|}{\tau}+\sqrt{\frac{L^{2}(m-1)+\frac{1}{\tau^{2}}}{\frac{1}{2\tau}-L\sqrt{m-1}}\cdot\Big{(}\mathcal{F}(\widetilde{\rho}^{k-1})-\mathcal{F}(\widetilde{\rho}^{k})+\frac{\|\varepsilon^{k}\|^{2}}{2\tau}\Big{)}}\bigg{]}.$
Here, we apply Lemma 7 and the fact that $\operatorname{W}_{2}(\widetilde{\rho}^{k},\rho^{\ast})\leq D_{\mathcal{X}}$. If $\mathcal{F}(\widetilde{\rho}^{k})-\mathcal{F}(\rho^{\ast})\geq\sqrt{3}D_{\mathcal{X}}\frac{\|\varepsilon^{k}\|}{\tau}$, we have
@@
Applying Lemma 15 with
@@ Appendix B Proof of Lemmas in the Main Theorems
To prove the two steps mentioned in the sketch at the beginning of Appendix A, the most important result is the following first-order optimality condition (FOC). The proof, which will be postponed to Appendix C.2, is similar to Proposition 8.7 in (Santambrogio, 2015) for $h(x)=x\log x$ and $W=0$.
Lemma 9 (FOC).
@@ B.1 Proof of Lemma 1
By Lemma 9, we have $T_{\rho_{j}^{k+1}}^{\rho_{j}^{k}}-\textrm{Id}=\tau\nabla V_{j}^{k}+\tau\nabla h_{j}^{\prime}(\rho_{j}^{k+1})+\tau\nabla_{\operatorname{W}_{2}}\mathcal{W}_{j}(\rho_{j}^{k+1})$. This implies
@@
$=\sum_{j=1}^{m}\int_{\mathcal{X}_{j}}\bigg{\langle}\frac{T_{\rho_{j}^{k+1}}^{\rho_{j}^{k}}-\textrm{Id}}{\tau}-\big{(}\nabla V_{j}^{k}+\nabla h_{j}^{\prime}(\rho_{j}^{k+1})+\nabla_{\operatorname{W}_{2}}\mathcal{W}_{j}(\rho_{j}^{k+1})\big{)},T_{\rho_{j}^{k+1}}^{\rho_{j}^{k}}-\textrm{Id}\bigg{\rangle}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\big{\langle}\nabla V_{j}^{k+1},T_{\rho_{j}^{k+1}}^{\rho_{j}^{k}}-\textrm{Id}\big{\rangle}\,{\rm d}\rho_{j}^{k+1}$
@@
$\qquad\qquad\qquad\qquad\qquad\qquad-\int_{\mathcal{X}_{j}}\Big{\langle}\nabla_{\operatorname{W}_{2}}\mathcal{W}_{j}(\rho_{j}^{k+1}),T_{\rho_{j}^{k+1}}^{\rho_{j}^{k}}-\textrm{Id}\Big{\rangle}\,{\rm d}\rho_{j}^{k+1}+\frac{1}{\tau}\operatorname{W}_{2}^{2}(\rho_{j}^{k+1},\rho_{j}^{k})\bigg{]}$
$\stackrel{{\scriptstyle(\textrm{i})}}{{\geq}}\sum_{j=1}^{m}\bigg{[}\int_{\mathcal{X}_{j}}\Big{\langle}\nabla V_{j}^{k+1}-\nabla V_{j}^{k},T_{\rho_{j}^{k+1}}^{\rho_{j}^{k}}-\textrm{Id}\Big{\rangle}\,{\rm d}\rho_{j}^{k+1}-\Big{(}\mathcal{H}_{j}(\rho_{j}^{k})-\mathcal{H}_{j}(\rho_{j}^{k+1})\Big{)}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\Big{(}\mathcal{W}_{j}(\rho_{j}^{k})-\mathcal{W}_{j}(\rho_{j}^{k+1})\Big{)}\bigg{]}+\frac{1}{\tau}\operatorname{W}_{2}^{2}(\rho^{k+1},\rho^{k})$
$\stackrel{{\scriptstyle(\textrm{ii})}}{{\geq}}-\sum_{j=1}^{m}\operatorname{W}_{2}(q_{j}^{k+1},q_{j}^{k})\cdot\Big{\|}\nabla V_{j}^{k+1}-\nabla V_{j}^{k}\Big{\|}_{L^{2}(\rho_{j}^{k+1};\mathcal{X}_{j})}-\Big{(}\mathcal{H}(\rho^{k})-\mathcal{H}(\rho^{k+1})\Big{)}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\Big{(}\mathcal{W}(\rho^{k})-\mathcal{W}(\rho^{k+1})\Big{)}+\frac{1}{\tau}\operatorname{W}_{2}^{2}(\rho^{k+1},\rho^{k})$
$\stackrel{{\scriptstyle(\textrm{iii})}}{{\geq}}-\sum_{j=1}^{m}\operatorname{W}_{2}(\rho_{j}^{k+1},\rho_{j}^{k})\cdot L\operatorname{W}_{2}(\rho_{-j}^{k+1},\rho_{-j}^{k})-\Big{(}\mathcal{H}(\rho^{k})-\mathcal{H}(\rho^{k+1})\Big{)}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\Big{(}\mathcal{W}(\rho^{k})-\mathcal{W}(\rho^{k+1})\Big{)}+\frac{1}{\tau}\operatorname{W}_{2}^{2}(\rho^{k+1},\rho^{k})$
$\stackrel{{\scriptstyle\textrm{(iv)}}}{{\geq}}-L\sum_{j=1}^{m}\bigg{[}\frac{\sqrt{m-1}}{2}\operatorname{W}_{2}^{2}(\rho_{j}^{k+1},\rho_{j}^{k})+\frac{1}{2\sqrt{m-1}}\operatorname{W}_{2}^{2}(\rho_{-j}^{k+1},\rho_{-j}^{k})\bigg{]}$
$\qquad\qquad\qquad\qquad\qquad-\Big{(}\mathcal{H}(\rho^{k})-\mathcal{H}(\rho^{k+1})\Big{)}-\Big{(}\mathcal{W}(\rho^{k})-\mathcal{W}(\rho^{k+1})\Big{)}+\frac{1}{\tau}\operatorname{W}_{2}^{2}(\rho^{k+1},\rho^{k})$
@@
Here, (i) is due to the fact that both $\mathcal{H}_{j}$ and $\mathcal{W}_{j}$ are convex along geodesics; (ii) is by Cauchy–Schwarz inequality; (iii) is by Lemma 12; (iv) is by AM-GM inequality.
@@
$\geq m\mathcal{V}(\rho^{k+1})+(m-1)\Big{[}\big{(}\tau^{-1}-L\sqrt{m-1}\big{)}\operatorname{W}_{2}^{2}(\rho^{k+1},\rho^{k})-\Big{(}\mathcal{H}(\rho^{k})-\mathcal{H}(\rho^{k+1})\Big{)}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\Big{(}\mathcal{W}(\rho^{k})-\mathcal{W}(\rho^{k+1})\Big{)}\Big{]}.$
@@ B.2 Proof of Lemma 2
Here, both (i) and (ii) are due to Cauchy–Schwarz inequality, and (iii) is by Lemma 10. By Assumption D ($\lambda$-QG condition), we have
@@
B.3 Proof of Lemma 3
Proof  By the definition of $\rho_{j}^{k}$, we have
@@ B.3 Proof of Lemma 3
B.4 Proof of Lemma 4
Proof  Let $Y_{j}=\sum_{l=0}^{M-1}I\{j_{l}=j\}$ and $A_{j}=\{Y_{j}=0\}\cup\{Y_{j}>T\}$ be a event. Then
@@ B.4 Proof of Lemma 4
Applying Chernoff’s inequality to Bernoulli distribution (Theorem 2.3.1 in(Vershynin, 2018)) yields
@@
we have $2m^{1-R}=2\delta$.
B.5 Proof of Lemma 5
Proof  By the definition of $\rho_{j_{l}}^{k,l+1}$, we have
@@ B.5 Proof of Lemma 5
Here, both (i) and (ii) are due to $\rho_{-j_{l}}^{k,l+1}=\rho_{-j_{l}}^{k,l}$.
B.6 Proof of Lemma 6
Proof  Similar to the argument in Lemma 2, we have
@@ B.7 Proof of Lemma 7
$\sum_{j=1}^{m}\mathcal{V}(\widetilde{\rho}_{j}^{k+1}\otimes\widetilde{\rho}_{-j}^{k})$ $\leq m\mathcal{V}(\widetilde{\rho}^{k})+\mathcal{H}(\widetilde{\rho}^{k})+\mathcal{W}(\widetilde{\rho}^{k})$
$\qquad\qquad\qquad-\mathcal{H}(\widetilde{\rho}^{k+1})-\mathcal{W}(\widetilde{\rho}^{k})-\frac{1}{2\tau}\operatorname{W}_{2}^{2}(\widetilde{\rho}^{k+1},\widetilde{\rho}^{k})+\frac{\|\varepsilon^{k+1}\|^{2}}{2\tau}.$
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\Big{\langle}\nabla V_{j}^{k+1},T_{\widetilde{\rho}_{j}^{k+1}}^{\widetilde{\rho}_{j}^{k}}-\textrm{Id}\Big{\rangle}\,{\rm d}\widetilde{\rho}_{j}^{k+1}$
@@
$\quad\qquad\qquad\qquad\qquad\qquad\quad-\int_{\mathcal{X}_{j}}\Big{\langle}\nabla_{\operatorname{W}_{2}}\mathcal{W}_{j}(\widetilde{\rho}_{j}^{k+1}),T_{\widetilde{\rho}_{j}^{k+1}}^{\widetilde{\rho}_{j}^{k}}-\textrm{Id}\Big{\rangle}\,{\rm d}\widetilde{\rho}_{j}^{k+1}+\frac{1}{\tau}\operatorname{W}_{2}^{2}(\widetilde{\rho}_{j}^{k+1},\widetilde{\rho}_{j}^{k})$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\frac{1}{\tau}\int_{\mathcal{X}_{j}}\Big{\langle}\eta_{j}^{k+1},T_{\widetilde{\rho}_{j}^{k+1}}^{\widetilde{\rho}_{j}^{k}}-\textrm{Id}\Big{\rangle}\,{\rm d}\widetilde{\rho}_{j}^{k+1}\bigg{]}$
@@
$\quad\qquad\qquad\qquad\quad-\int_{\mathcal{X}_{j}}\Big{\langle}\nabla_{\operatorname{W}_{2}}\mathcal{W}_{j}(\widetilde{\rho}_{j}^{k+1}),T_{\widetilde{\rho}_{j}^{k+1}}^{\widetilde{\rho}_{j}^{k}}-\textrm{Id}\Big{\rangle}\,{\rm d}\widetilde{\rho}_{j}^{k+1}+\frac{1}{2\tau}\operatorname{W}_{2}^{2}(\widetilde{\rho}_{j}^{k+1},\widetilde{\rho}_{j}^{k})-\frac{(\varepsilon_{j}^{k+1})^{2}}{2\tau}\bigg{]}$
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\big{[}\mathcal{W}(\widetilde{\rho}^{k})-\mathcal{W}(\widetilde{\rho}^{k+1})\big{]}+\frac{1}{2\tau}\operatorname{W}_{2}^{2}(\widetilde{\rho}^{k+1},\widetilde{\rho}^{k})-\frac{\|\varepsilon^{k+1}\|^{2}}{2\tau}$
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\big{[}\mathcal{W}(\widetilde{\rho}^{k})-\mathcal{W}(\widetilde{\rho}^{k+1})\big{]}+\frac{1}{2\tau}\operatorname{W}_{2}^{2}(\widetilde{\rho}^{k+1},\widetilde{\rho}^{k})-\frac{\|\varepsilon^{k+1}\|^{2}}{2\tau}$
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\big{[}\mathcal{W}(\widetilde{\rho}^{k})-\mathcal{W}(\widetilde{\rho}^{k+1})\big{]}+\frac{1}{2\tau}\operatorname{W}_{2}^{2}(\widetilde{\rho}^{k+1},\widetilde{\rho}^{k})-\frac{\|\varepsilon^{k+1}\|^{2}}{2\tau}$